OpenStudy (anonymous):

Help Please! Reduce rational expression, x to the 2 power-5x+6 x-4 _________________________ . ________________ x to the 2 power-4x x-3 Thank you.

OpenStudy (anonymous):

Try the same approach I helped with earlier. x^2 - 4x can be simplified. So, can x^2-5x+6 be by factoring. Then, a simple matter of canceling out common terms.

OpenStudy (anonymous):

i got x+5 over x-2 ? sorry, i'm having a hard time with this.....

OpenStudy (anonymous):

x^2 - 4x = x(x-4) x^2 - 5x + 6 = (x-3)(x-2) (x-3)(x-2)(x-4)/x(x-4)(x-3) = x-2/x